Bill's upstream canoe trip took 1.5 hours longer than the return trip. If he averaged 2 mph on the upstream trip and 3 mph on the return trip, then how long was the upstream trip?
Answer by mananth(16949)

Bill's upstream canoe trip took 1.5 hours longer than the return trip. If he averaged 2 mph on the upstream trip and 3 mph on the return trip, then how long was the upstream trip?
let the return trip take x hours
upstream will take x+1.5 hours
distance upstream = 2(x+1.5)
distance down stream = 3x
the distance is the same both ways
2(x+1.5)= 3x
2x+3=3x
3x-2x=3
x=3 down stream time
up stream will be 3+1.5 hours =4.5 hours